# Global convergence of a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems

### v1

@@
Abstract: We propose a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems (MFQNMO). The method employs a common BFGS matrix to approximate the Hessian matrix of all objective functions in each iteration. This matrix is updated using function and gradient information from the previous step. Our analysis confirms the convergence of the method without relying on convexity assumptions. We also establish a local superlinear convergence rate for MFQNMO under mild conditions and validate its effectiveness through experiments on both nonconvex and convex test problems.
Keywords: Multiobjective optimization, quasi-Newton methods, nonconvex, Wolfe line search, global convergence.
@@ 1 Introduction
In this paper, we consider the following unconstrained multiobjective optimization problems (MOP):
$\min_{x\in R^{n}}f(x)=\left(f_{1}\left(x\right),f_{2}\left(x\right),\ldots,f_{m}\left(x\right)\right)^{T},$ (MOP)
@@
where $f_{i}\left(x\right){:}R^{n}\to R$ is continuously differentiable function, for any $i\in[m]$. Multiobjective optimization is widely used in many research areas, which can be found in engineering [1, 2], environmental analysis [3, 4], machine learning [5, 6], etc.
@@
In 2000, Fliege and Svaiter proposed the steepest descent methods for multiobjective optimization [7]. Building on their pioneering work, researchers have conducted extensive studies on the first-order descent algorithm for solving (MOP), (see e.g. [11, 12, 13, 14]). The first-order algorithm is often favored for its lower cost compared to the second-order method, as it avoids the computational burden of determining the second-order Hessian matrix. However, the trade-off is a typically slower convergence rate, particularly evident in scenarios involving ill-conditioned problems.
@@
The ill-conditioned problem may exhibit large curvature variations at certain points, leading to highly unstable curvature values. The Newton method [8] is effective in such cases as it captures the local curvature of functions and adjusts the direction more efficiently, leading to faster rate compared to first-order algorithms. Fliege et al. [8] proved the superlinear convergence rate of the Newton method. However, despite its efficiency, the computational cost per iteration is notably high, particularly for large-scale problems.
@@
To balance the computational cost and the effectiveness, the quasi-Newton method is considered. Different from single objective optimization, the quasi-Newton method [9] uses different matrices to approximate the Hessian matrix of individual objective functions in each iteration to capture the curvature information. Numerous variants of the quasi-Newton method have been proposed. Lai et al. [16] proposed a quasi-Newton method for constructing approximate matrices using q-derivate. In [17], the global convergence of the quasi-Newton method is established by the average nonmonotone line search. Morovati et al. [18] extended both the self-scaling BFGS (SS-BFGS) method and the Huang BFGS (H-BGFS) method for multiobjective optimization problems. Prudente and Souza [19], operating under the assumption of no convexity, proposed a modified BFGS formula. They demonstrated that their algorithm is well-defined under the Wolf line search without convexity assumption. Different from the case of single-objective, the computational cost per iteration still be high when the number of objective functions increases for multiobjective optimization problems. To address this, Ansary and Panda [15] utilized a common matrix to approximate all Hessian matrices in each iteration. Lapucci and Mansueto [21] extended the limited memory Quasi-Newton approach for MOPs (L-BFGS). However, all the improved algorithms mentioned above require convexity conditions to ensure global convergence. In [10, 20], the global convergence can be guaranteed with Wolf line search in the nonconvex case. The approximate Secant equation in reference [20] depends on some parameters, and updating the BFGS formula under strong convexity conditions does not always ensure a satisfactory approximation of the Secant equation.
@@
For MOPs, there are relatively few studies on quasi-Newton methods in nonconvex cases. This raises the question of whether alternative methods exist that can balance between algorithmic convergence and per-iteration cost for nonconvex MOPs. In this paper, inspired by [20, 15], we proposed a modified BFGS-type method based on function information for nonconvex MOPs (MFQNMO). By utilizing both function and gradient information from the previous iterations, we construct a common modified BFGS-type formula to approximate the Hessian matrices of all objective functions in each iteration. We verify that the modified BGFS-type matrix is positive definite for each iteration and establish a global convergence for MFQNMO without convexity assumption.
@@
The structure of this paper is as follows: In sect. 2, we give some notations and definitions for our later use. Sect. 3 recalls several multiobjective gradient descent methods. In sect. 4, we introduce the proposed MFQNMO method and establish its convergence analysis. Numerical results are presented in sect. 5. At the end of the paper, we draw some conclusions.
@@ 2 Preliminaries
We give some notations used in this paper.
@@
$\bullet$ ${{\Delta}_{m}}=\left\{\lambda:\sum\limits_{i\in\left[m\right]}{{{\lambda}_{i}}=1},{{\lambda}_{i}}\geq 0,\ i\in\left[m\right]\right\}.$
@@
$\bullet$ ${}Jf\left(x\right)\in{{R}^{m\times n}}$, $\nabla{{f}_{i}}\left(x\right)\in{{R}^{n}}$ ,$\nabla^{2}f_{i}(x)$ the Jacobian matrix, the gradient of ${{f}_{i}}$ and the Hessian matrix at $x$, respectively.
$\bullet$ $\mathcal{D}(x,d)\coloneqq\max\limits_{i\in[m]}\nabla f_{i}(x)^{\top}d.$
$\bullet$ For all $B\in R^{n\times n}$, $B\succ 0$ is equivalent to $B$ is positive definite.
$\bullet$ For $x,y\in R^{n}$, $\langle x,y\rangle_{B}:=x^{\top}By$ and $\|x\|_{B}:=\sqrt{\langle x,x\rangle_{B}}$.
We define partial order for our later use:
$A\preceq B\Leftrightarrow B-A\succeq 0,\quad A,B\in R^{n\times n}$
$f\left(y\right)\leq f\left(x\right)\Leftrightarrow{{f}_{i}}\left(x\right)-{{f}_{i}}\left(y\right)\geq 0,\ i\in\left[m\right].$
@@
[7] A point ${{x}^{*}}\in{{R}^{n}}$ is called the Pareto (weakly Pareto ) solution for MOP, if there exists no point $x\in{{R}^{n}}$ such that $f\left(x\right)\leq f\left({{x}^{*}}\right)$ and $f\left(x\right)\neq f\left({{x}^{*}}\right)~{}\left(f\left(x\right)<f\left({{x}^{*}}\right)\right)$.
@@
[7] A point ${{x}^{*}}\in{{R}^{n}}$ is called a Pareto critical point for MOP, if exists $i\in\left[m\right]$ such that $\left\langle\nabla{{f}_{i}}\left({{x}^{*}}\right),d\right\rangle\geq 0,$ for all $d\in{{R}^{n}}.$
@@
[7] A vector $d\in{{R}^{n}}$ is called descent direction for $f$ at ${x}^{*}$, if $\left\langle\nabla{{f}_{i}}\left({{x}^{*}}\right),d\right\rangle<0,$ for all $i\in\left[m\right]$.
@@
3 Gradient descent methods for MOPs.
@@
For a given $x\in{{R}^{n}}$, consider the following subproblem (3.1):
$\underset{d\in{{R}^{n}}}{\mathop{\min}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}.$ (3.1)
@@
$d\left(x\right)=\underset{d\in{{R}^{n}}}{\mathop{\text{argmin}}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}},$
@@
$\theta\left(x\right)=\underset{d\in{{R}^{n}}}{\mathop{\min}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}.$
@@
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t+\frac{1}{2}\|d\|^{2}\\ \text{s.t.}&\langle\nabla f_{i}(x),d\rangle\leq t,\quad i\in[m].\end{array}$ (3.2)
According to the Slater condition, strong duality is established. The dual problem has the same minimum value as (3.2). We consider that the dual problem of problem (3.2):
$\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\max}}\,\underset{\left(t,d\right)\in R\times{{R}^{n}}}{\mathop{\min}}\,L\left(\lambda,\left(t,d\right)\right),$
@@
where $\lambda_{SD}\left(x\right)={{\left({{\lambda}_{1}^{SD}}\left(x\right),{{\lambda}_{2}^{SD}}\left(x\right),\ldots,{{\lambda}_{m}^{SD}}\left(x\right)\right)}^{T}}$is the solution of the following dual problem
@@
$\theta_{SD}\left(x\right)=-\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}^{SD}}\left(x\right)\nabla{{f}_{i}}\left(x\right)\right\|}^{2}}=-\frac{1}{2}{{\left\|d_{SD}\left(x\right)\right\|}^{2}}\leq 0,$
$\left\langle\nabla{{f}_{i}}\left(x\right),d_{SD}\left(x\right)\right\rangle\leq-{{\left\|d_{SD}\left(x\right)\right\|}^{2}}\leq 0.$
@@
$(i)$ The point $x\in{{R}^{n}}$ is not Pareto critical;
@@
$(b)$ The function $d_{SD}\left(x\right)$ and $\theta_{SD}\left(x\right)$ are continuous, for all $x\in{{R}^{n}}.$
@@
$\min_{d\in R^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d,$ (3.4)
where $B_{i}(x)\in R^{n\times n}$, approximates the Hessian matrix $\nabla^{2}f_{i}(x)$ and $B_{i}(x)\succ 0$, for $i\in[m]$. $\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d$ is strongly convex, for $i\in[m]$. The subproblem 3.4 has a unique minimizer. We denote that $\theta_{q}\left(x\right)$ and $d_{q}\left(x\right)$ be the optimal value and solution in problem (3.4), respectively,
$d_{q}(x)=\mathop{\text{argmin}}_{d\in R^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d,$
@@
$\theta_{q}(x)=\max_{i\in[m]}\nabla f_{i}(x)^{\top}d(x)+\frac{1}{2}d(x)^{\top}B_{i}(x)d(x).$
@@
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t\\ \text{s.t.}&\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d\leq t,\quad\forall i\in[m].\end{array}$
@@
$d_{q}(x)=-\left[\sum_{i=1}^{m}\lambda_{i}^{q}(x)B_{i}(x)\right]^{-1}(\sum_{i=1}^{m}\lambda_{i}^{q}(x)\nabla f_{i}(x)),$ (3.5)
where $\lambda_{q}\left(x\right)={{\left({{\lambda}_{1}^{q}}\left(x\right),{{\lambda}_{2}^{q}}\left(x\right),\ldots,{{\lambda}_{m}^{q}}\left(x\right)\right)}^{T}}$is the solution of the following dual problem
$-\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\min}}\,\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}}\nabla{{f}_{i}}\left(x\right)\right\|^{2}_{\left[\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}{B_{i}(x)}}\right]^{-1}}}}.$
@@
Even if quasi-Newton avoids calculating the second-order Hessian matrix, the per-iteration cost is still expensive. Recently, Ansary and Panda [15] proposed a modified quasi-Newton method (MQNMO) that utilized a single positive matrix to approximate all the Hessian matrices. The Modified quasi-Newton direction $d(x)$ is the optimal solution to the following subproblem 3.6:
$\min_{d\in R^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B(x)d,$ (3.6)
where $B(x)\in R^{n\times n}$ and $B(x)\succ 0$. Problem (3.6) is equivalent to the convex quadratic optimization problem as follows
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t\\ \text{s.t.}&\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B(x)d\leq t,\quad\forall i\in[m].\end{array}$
According to the KKT condition, we can obtain the modified quasi-Newton direction
@@
where $\lambda\left(x\right)={{\left({{\lambda}_{1}}\left(x\right),{{\lambda}_{2}}\left(x\right),\ldots,{{\lambda}_{m}}\left(x\right)\right)}^{T}}$is the solution of the following dual problem (3.9)
@@
where $s_{k}=x_{k+1}-x_{k}$ and $y_{k}=\sum_{i=1}^{m}\lambda_{i}^{k}(\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})).$
4 MFQNMO:Modified BFGS method based on function information for MOPs.
In this section, we propose a modified BFGS method based on function information for MOPs and analyze its properties. The formal description of the algorithm is as follows.
@@
1:Choose starting point $x^{0}\in\mathbb{R}^{n}$ and parameters $\sigma_{1},\sigma_{2}\in(0,1).$ Let $k=0$ and $B_{0}\succ 0.$
2:Compute the solution ${{\lambda}_{k}}={{\left({{\lambda}_{1}}\left({{x}_{k}}\right),{{\lambda}_{2}}\left({{x}_{k}}\right),\ldots,{{\lambda}_{m}}\left({{x}_{k}}\right)\right)}^{T}}$ of problem (3.9) and ${{d}_{k}}$, $\theta_{k}$ as (3.7), (3.8).
@@
$F_{i}(x_{k}+\alpha_{k}d_{k})$ $\leq F_{i}(x_{k})+\rho\alpha_{k}\mathcal{D}(x_{k},d_{k}),\quad\forall i\in[m],$ (4.1)
$\mathcal{D}(x^{k}+\alpha_{k}d_{k},d_{k})$ $\geq\sigma\mathcal{D}(x_{k},d_{k}),$ (4.2)
@@
where $s_{k}=x_{k+1}-x_{k}$ , $\mu_{i}^{k}=\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})$ and $y_{k}=\sum_{i=1}^{m}\lambda_{i}^{k}\mu_{i}^{k}$.
@@
The BFGS matrix (4.6) updated by Algorithm 1 is positive definite.
@@
At the k-th iteration, we only need to consider the case when $x_{k}$ is not the Pareto critical point. By the definition of $\gamma_{k}$, $\eta_{k}$ and $m_{k}$, we have
@@
$\geq\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x_{k})-f_{i}(x_{k+1})\right)\|s_{k}\|^{2}$ (4.9)
@@
Then for all nonzero vector $x\in R^{n}$,

### v2

@@
Abstract: In this paper, based on function information, we propose a modified BFGS-type method for nonconvex multiobjective optimization problems (MFQNMO). In the multiobjective quasi-Newton method (QNMO), each iteration involves separately approximating the Hessian matrix for each component objective function, which results in significant storage and computational burdens. MFQNMO employs a common BFGS-type matrix to approximate the Hessian matrix of all objective functions in each iteration. This matrix is updated using function information from the previous step. This approach strikes a balance between efficacy and computational cost. We confirm the convergence of the method without relying on convexity assumptions, under mild conditions, we establish a local superlinear convergence rate for MFQNMO. Furthermore, we validate its effectiveness through experiments on both nonconvex and convex test problems.
Keywords: Nonconvex multiobjective optimization, Quasi-Newton methods, Wolfe line search, Global convergence, Superlinear convergence.
@@ 1 Introduction
In this paper, we consider the following unconstrained multiobjective optimization problem (MOP):
$\min_{x\in\mathbb{R}^{n}}f(x)=\left(f_{1}\left(x\right),f_{2}\left(x\right),\ldots,f_{m}\left(x\right)\right)^{\top},$ (MOP)
@@
where $f_{i}\left(x\right){:}\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable function, for any $i\in[m]$. Multiobjective optimization is widely used in many research areas, which can be found in engineering [23, 18], environmental analysis [13, 1], machine learning [32, 28], etc.
@@
In 2000, Fliege and Svaiter [7] proposed the steepest descent method for multiobjective optimization (SDMO). Building on their pioneering work, researchers have conducted extensive studies on the first-order descent algorithm for solving (MOP), (see e.g. [31, 19, 2, 15]). The first-order algorithm is often favored for its lower cost than the second-order method, as it avoids the computational burden of determining the Hessian matrix. However, this advantage comes with a trade-off: first-order methods typically exhibit slower convergence rates, especially ill-conditioned problems.
@@
Ill-conditioned problems exhibit large curvature variations at certain points, resulting in highly unstable curvature values. The Newton method [8] is effective in such cases as it captures the local curvature of functions and adjusts the direction more efficiently, leading to a faster rate compared to first-order methods. Fliege et al. [8] proved the superlinear convergence rate of the Newton method. However, despite its efficiency, the computational cost per iteration is significantly higher, particularly for large-scale problems.
@@
To reduce the computational costs incurred by the direct calculation of the Hessian matrix, the quasi-Newton method is considered. Unlike single objective optimization, for individual objective functions, the quasi-Newton methods [24, 27] employ different matrices to approximate their Hessian matrices, capturing the local curvature. Numerous variants of the quasi-Newton method have been proposed for multi-objective problems (MOPs). In [20], the global convergence of the quasi-Newton method is established under average nonmonotone line search. Morovati et al. [21] extended both the self-scaling BFGS (SS-BFGS) method and the Huang BFGS (H-BGFS) method for multiobjective optimization problems. Prudente and Souza [26] proposed a modified BFGS update formula under nonconvex assumptions, ensuring that the algorithm is well-defined. Furthermore, under the assumption of strong convexity, they establish the global convergence and local superlinear convergence rate of the method with Wolfe line search.
In large-scale multiobjective optimization problems, an increase in the number of objective functions results in an exponential growth in the scale of the matrices that need to be stored and computed, particularly in cases with high-dimensional variable spaces. Consequently, it is impractical to calculate or approximate the Hessian matrix for each objective function separately in large-scale scenarios. To address this, Ansary and Panda [3] utilized a common matrix to approximate all Hessian matrices in each iteration and establish the global convergence under an aggregated Wolf line search. Lapucci and Mansueto [14] extended the limited memory Quasi-Newton approach (L-BFGS) for large-scale MOPs, and the global convergence with R-linear convergence rate is obtained under the strong convexity assumption. However, all the improved algorithms mentioned above require convexity conditions to ensure global convergence. Chen et al. [5] proposed a variable metric method for unconstrained MOPs. Based on an assumption weaker than the convexity condition, this method establishes strong convergence. Additionally, they mentioned that solving the quadratic constrained subproblem of the second-order method is more difficult than solving the direction-finding dual problem of the first-order method. In [25], the global convergence can be guaranteed with Wolf line search in the nonconvex case.
@@
For MOPs, there are relatively few studies on quasi-Newton methods in nonconvex cases. This raises the question of whether alternative quasi-Newton method exists that can strike a balance between global convergence and per-iteration cost for nonconvex MOPs. In this paper, inspired by [25, 3], by utilizing function information, we construct a common modified BFGS-type matrix to approximate the Hessian matrices of all objective functions in each iteration. We proposed a modified BFGS-type method based on function information for nonconvex MOPs (MFQNMO). We verify that the modified BGFS-type matrix is positive definite for each iteration and establish a global convergence for MFQNMO without convexity assumption. Under some mild conditions, we state the superlinear convergence property of MFQNMO. It also shows superior performance in numerical experiments. Our work has the following advantages:
$\bullet$ Rather than approximating the Hessian matrix for each objective function individually, our method utilizes a single common matrix to capture the curvature information. This approach strategically balances the effectiveness of the optimization algorithm with its computational cost.
$\bullet$ For nonconvex functions, most versions of quasi-Newton methods encounter difficulties in achieving convergence. Therefore, the MFQNMO method proposed in this paper for nonconvex unconstrained problems is particularly significant and valuable.
$\bullet$ our method’s update process for the BFGS-type formula does not rely on any specific parameter selections. Instead, it only utilizes information from prior iterations.
@@
The structure of this paper is as follows: In Sect. 2, we give some notations and definitions for our later use. Sect. 3 recalls several multiobjective descent methods, including SDMO, QNMO and MQNMO. In Sect. 4, we introduce the proposed MFQNMO method and establish its convergence analysis. Numerical results are presented in Sect. 5. At the end of the paper, we draw some conclusions.
@@ 2 Preliminaries
In this section, we recall some notations, definitions and results used in the sequel. Let $\mathbb{R}^{n}$ be the $n$-dimensional Euclidean space, and $\mathbb{R}_{+}^{n}$ denote nonnegative orthant of $\mathbb{R}^{n}$. $N$ represents the set of natural numbers. For any $x:=(x_{1},\ldots,x_{n})$ and $y:=(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}$, we define:
$x\geqslant y$$\Leftrightarrow$ $x-y\in\mathbb{R}_{+}^{n}\setminus\{0\};$
$x\ngeqslant y$$\Leftrightarrow$ $x-y\notin\mathbb{R}_{+}^{n}\setminus\{0\};$
$x>y$$\Leftrightarrow$ $x-y\in\texttt{int}\mathbb{R}_{+}^{n};$
$x\not>y$$\Leftrightarrow$ $x-y\notin\texttt{int}\mathbb{R}_{+}^{n}.$
The following are some notations that will be used.
@@
$\bullet$ ${{\Delta}_{m}}=\left\{\lambda:\sum\limits_{i\in\left[m\right]}{{{\lambda}_{i}}=1},{{\lambda}_{i}}\geqslant 0,\ i\in\left[m\right]\right\}.$
@@
$\bullet$ ${}Jf\left(x\right)\in{{\mathbb{R}}^{m\times n}}$, $\nabla{{f}_{i}}\left(x\right)\in{{\mathbb{R}}^{n}}$ ,$\nabla^{2}f_{i}(x)$ the Jacobian matrix, the gradient of ${{f}_{i}}$ and the Hessian matrix at $x$, respectively.
$\bullet$ $\mathcal{D}(x,d)\coloneqq\max\limits_{i\in[m]}\nabla f_{i}(x)^{\top}d$, $d\in\mathbb{R}^{n}.$
$\bullet$ For all $B\in\mathbb{R}^{n\times n}$, $B\succ 0$ means that $B$ is positive definite, $B\succeq 0$ means that $B$ is positive semi-definite.
$\bullet$ For $x,y\in\mathbb{R}^{n}$, $\langle x,y\rangle_{B}:=x^{\top}By$ and $\|x\|_{B}:=\sqrt{\langle x,x\rangle_{B}}$.
$\bullet$ For $x\in\mathbb{R}$, $\lceil x\rceil$ is the least integer greater than or equal to x.
$\bullet$ $A\preceq B\Leftrightarrow B-A\succeq 0,\quad A,B\in\mathbb{R}^{n\times n}$
@@
[7] A point ${{x}^{*}}\in{{\mathbb{R}}^{n}}$ is called the Pareto (weakly Pareto ) solution for MOP, if there exists no point $x\in{{\mathbb{R}}^{n}}$ such that $f\left(x\right)\leqslant f\left({{x}^{*}}\right)$ and $f\left(x\right)\neq f\left({{x}^{*}}\right)~{}\left(f\left(x\right)<f\left({{x}^{*}}\right)\right)$.
@@
[7] A point ${{x}^{*}}\in{{\mathbb{R}}^{n}}$ is called a Pareto critical point for MOP, if exists $i\in\left[m\right]$ such that $\left\langle\nabla{{f}_{i}}\left({{x}^{*}}\right),d\right\rangle\geqslant 0,$ for all $d\in{{\mathbb{R}}^{n}}.$
@@
[7] A vector $d\in{{\mathbb{R}}^{n}}$ is called descent direction for $f$ at ${x}^{*}$, if $\left\langle\nabla{{f}_{i}}\left({{x}^{*}}\right),d\right\rangle<0,$ for all $i\in\left[m\right]$.
@@
3 Descent methods for MOPs.
@@
For a given $x\in{{\mathbb{R}}^{n}}$, consider the following subproblem (3.1):
$\underset{d\in{{\mathbb{R}}^{n}}}{\mathop{\min}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}.$ (3.1)
@@
$d\left(x\right)=\underset{d\in{{\mathbb{R}}^{n}}}{\mathop{\text{argmin}}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}},$
@@
$\theta\left(x\right)=\underset{d\in{{\mathbb{R}}^{n}}}{\mathop{\min}}\,\underset{i\in[m]}{\mathop{\max}}\,\left\langle\nabla{{f}_{i}}\left(x\right),d\right\rangle+\frac{1}{2}{{\left\|d\right\|}^{2}}.$
@@
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t+\frac{1}{2}\|d\|^{2}\\ \text{s.t.}&\langle\nabla f_{i}(x),d\rangle\leqslant t,\quad i\in[m].\end{array}$ (3.2)
We consider that the dual problem of problem (3.2):
$\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\max}}\,\underset{\left(t,d\right)\in\mathbb{R}\times{{\mathbb{R}}^{n}}}{\mathop{\min}}\,L\left(\lambda,\left(t,d\right)\right),$
@@
where $\lambda_{SD}\left(x\right)={{\left({{\lambda}_{1}^{SD}}\left(x\right),{{\lambda}_{2}^{SD}}\left(x\right),\cdots,{{\lambda}_{m}^{SD}}\left(x\right)\right)}^{\top}}$is the solution of the following dual problem
@@
$\theta_{SD}\left(x\right)=-\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}^{SD}}\left(x\right)\nabla{{f}_{i}}\left(x\right)\right\|}^{2}}=-\frac{1}{2}{{\left\|d_{SD}\left(x\right)\right\|}^{2}}\leqslant 0,$
$\left\langle\nabla{{f}_{i}}\left(x\right),d_{SD}\left(x\right)\right\rangle\leqslant-{{\left\|d_{SD}\left(x\right)\right\|}^{2}}\leqslant 0.$
@@
$(i)$ The point $x\in{{\mathbb{R}}^{n}}$ is not Pareto critical;
@@
$(b)$ The function $d_{SD}\left(x\right)$ and $\theta_{SD}\left(x\right)$ are continuous, for all $x\in{{\mathbb{R}}^{n}}.$
@@
$\min_{d\in\mathbb{R}^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d,$ (3.4)
where $B_{i}(x)\in\mathbb{R}^{n\times n}$, approximates the Hessian matrix $\nabla^{2}f_{i}(x).$ For $i\in[m]$, $B_{i}(x)\succ 0$ such that $\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d$ is strongly convex. Then, the subproblem 3.4 has a unique minimizer. We denote that $\theta_{q}\left(x\right)$ and $d_{q}\left(x\right)$ be the optimal value and solution in problem (3.4), respectively,
$d_{q}(x)=\mathop{\text{argmin}}_{d\in\mathbb{R}^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d,$
@@
$\theta_{q}(x)=\mathop{\text{min}}_{d\in\mathbb{R}^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d.$
@@
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t\\ \text{s.t.}&\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B_{i}(x)d\leqslant t,\quad\forall i\in[m].\end{array}$
@@
$d_{q}(x)=-\left[\sum_{i=1}^{m}\lambda_{i}^{q}(x)B_{i}(x)\right]^{-1}\left(\sum_{i=1}^{m}\lambda_{i}^{q}(x)\nabla f_{i}(x)\right),$ (3.5)
where $\lambda_{q}\left(x\right)={{\left({{\lambda}_{1}^{q}}\left(x\right),{{\lambda}_{2}^{q}}\left(x\right),\cdots,{{\lambda}_{m}^{q}}\left(x\right)\right)}^{\top}}$is the solution of the following dual problem
$-\underset{\lambda\in{{\Delta}_{m}}}{\mathop{\min}}\,\frac{1}{2}{{\left\|\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}}\nabla{{f}_{i}}\left(x\right)\right\|^{2}_{\left[\underset{i\in\left[m\right]}{\mathop{\sum}}\,{{\lambda}_{i}{B_{i}(x)}}\right]^{-1}}}},$
@@
Even if quasi-Newton avoids calculating the second-order Hessian matrix, the per-iteration cost is still expensive. Recently, Ansary and Panda [3] proposed a modified quasi-Newton method (MQNMO) that utilized a single positive matrix to approximate all the Hessian matrices. The Modified quasi-Newton direction $d(x)$ is the optimal solution to the following subproblem:
$\min_{d\in\mathbb{R}^{n}}\max_{i\in[m]}\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B(x)d,$ (3.6)
where $B(x)\in\mathbb{R}^{n\times n}$ and $B(x)\succ 0$. Problem (3.6) is equivalent to the convex quadratic optimization problem as follows
$\begin{array}[]{cl}\underset{(t,d)\in\mathbb{R}\times\mathbb{R}^{n}}{\text{min}}&\;t\\ \text{s.t.}&\nabla f_{i}(x)^{\top}d+\frac{1}{2}d^{\top}B(x)d\leqslant t,\quad\forall i\in[m].\end{array}$
Similar to the analysis in the previous section, we can obtain the modified quasi-Newton direction using the KKT condition
@@
where $\lambda\left(x\right)={{\left({{\lambda}_{1}}\left(x\right),{{\lambda}_{2}}\left(x\right),\cdots,{{\lambda}_{m}}\left(x\right)\right)}^{\top}}$is the solution of the following dual problem (3.9)
@@
where $s_{k}=x_{k+1}-x_{k}$ and $y_{k}=\sum\limits_{i=1}^{m}\lambda_{i}^{k}(\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})).$
4 MFQNMO:Modified BFGS-type method based on function information for MOPs.
In this section, we propose a modified BFGS-type method based on function information for MOPs and analyze its properties. The formal description of the algorithm is as follows.
@@
1:Choose starting point $x^{0}\in\mathbb{R}^{n}$. Let $10<\sigma_{1}<\sigma_{2}<$ 1, $k=0$, $\varepsilon\geqslant 0$ and $B_{0}\succ 0.$
2:Compute the solution ${{\lambda}_{k}}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\cdots,\lambda_{m}^{k}\right)}^{\top}}$ of problem (3.9) and ${{d}_{k}}$, $\theta_{k}$ as (3.7), (3.8).
@@
$f_{i}(x_{k}+\alpha_{k}d_{k})$ $\leqslant f_{i}(x_{k})+\sigma_{1}\alpha_{k}\mathcal{D}(x_{k},d_{k}),\quad\forall i\in[m],$ (4.1)
$\mathcal{D}(x^{k}+\alpha_{k}d_{k},d_{k})$ $\geqslant\sigma_{2}\mathcal{D}(x_{k},d_{k}),$ (4.2)
@@
where $s_{k}=x_{k+1}-x_{k}$ , $\mu_{i}^{k}=\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})$ and $y_{k}=\sum\limits_{i=1}^{m}\lambda_{i}^{k}\mu_{i}^{k}$.
@@
$x_{k}$ is not the Pareto critical point, then $B_{k}$ updated by Algorithm 1 is positive definite, for all $k\geq 0$.
@@
By the definition of $\gamma_{k}$, $\eta_{k}$ and $m_{k}$, we have
@@
$\geqslant\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x_{k})-f_{i}(x_{k+1})\right)\|s_{k}\|^{2}$ (4.9)
@@
Then for all vector $x\in\mathbb{R}^{n}$,